[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Sep 08, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Sep 8 21:34:26 UTC 2017


HSU Shift Log for Fri Sep 08 2017 17:34:18 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Monica and Chris Sisko were very helpful in ironing  
out some kinks in ATCF/NMAP2
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170908/59a53791/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list